RC Meeting and Run Plan, Wednesday, December 11th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • 5-pass setup took all of SWING and OWL
  • 5-pass beam-tuning ongoing
    • Working on beam centering with double carbon hole
    • On schedule to start Moller commissioning anytime
  • Both spectrometers at 30 degrees/2GeV; SHMS configured with defocussed optics (Q2 set +20%)[1]

Current Run Plan

  • Daily Run Plan and Run Status
  • Priorities
    • First: Spin Dance until done (tomorrow)
    • Next: commission and debug target polarity to be ready for 1-pass on Saturday DAY (AFTER Spin Dance)
    • Commission Moller (Dave Gaskell)
    • Run de-focused magnet tune for SHMS calorimeter calibration
    • Do BCM calibration if stable beam current >= 35uA available. You can find detailed instructions here
    • Do first reference cell N2 measurements if extra time
  • Tomorrow
    • Spin Dance during DAY (Dave Gaskell)
    • Reference cell N2 program
    • Be permissive of access as needed to debug target

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • There is no lock available for the ramp rate. This will be fixed Friday morning. Until then, request a more conservative 0.5uA/s ramp rate and carefully monitor.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.

Current Issues

  • Raster "breathing" visible on BPM [2] (not really an issue, just a set of interesting plots).
  • Unable to observe target polarization with NMR last night https://logbooks.jlab.org/entry/3750996]
    • Maybe caused by incorrect state of half wave plate, longitudinal laser polarization or incorrect field direction
  • EPR
    • Software operational [3]
    • EPR setup still needs debugging [4]; Suspected solution requires access.
  • Replaced "molten" heater fuse last night [5], Todd recommends using additional fans to prevent this from happening [6].

Resolved Issues

  • New alarms:
    • on the HMS polarity using the Hall Probe readout [7]
    • on the status of the insertable half-wave plate [8]
  • SHMS PID trigger tresholds as determined by Simona in Summer were found to be working perfectly for this run as well [9]

Planned Access to the Hall

  • Re-configure carbon foil target Friday DAY during RF recovery [10]
    • Needs coordination with RadCon and Techs. Keep in mind Techs only work to noon on Friday. Target foils may require RadWorker 2 (hopefully they won't). Need to start as early as possible.
  • Work on Laser RTD #4 (Murchhana & Mingyu, ~1h)
  • Debug EPR setup (Melanie)
  • Check coil polarities
  • Replace bunker fan
  • Check half-way plate status in laser box

Planned Accelerator Configuration Changes

  • Pass-change to 1-pass on Saturday 12/14 DAY
  • Pass-change to 5-pass on Tuesday 12/17 DAY
  • End-of-run (Holiday Break) Friday 12/20
  • Re-startup on 01/10/2020

Special Requests

  • Please be careful with our target :-)